SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

The Trigger Operation Meeting (May 1st, 2007)

DateTuesday, May 1st, 2007
Time2:00pm - 3:30pm PST
SLAC RoomMadrone conf room, ROB
Dial-in1-510-665-5437, pass code 6649#

Attendees: Kim Hojeong, Arik Kreisel, Selina Li, Rahmat Rahmat, Su Dong.

Agenda and Minutes

Operations

EMT

  • We need to test two spare boards with at least DCH and EMC, and make them good spare in IR2. Mark should leave an instruction in case the test can not be done before he leaves. The two repaired board should be kept in the teststand.

DCT/DCZ

  • Su Dong replaced the 3rd crate, which was broken, in the teststand with a proper crate on Friday. Now it is a little bit less convenient to access from side to the crate, but other than that it's working ok. Fang can test it.
  • Did Debbie finish DCC crate calibration?

GLT

  • Su Dong: compare distribution between old and new library. E, Y, G(*) objects have too small statistics to see in cosmics run, thus ignore any discrepancy in them. Others show good agreement. Conclude that the validation is ok, we'll change to the new library.
  • We need to test the GLT board with new library in, with all detectors up and running, but without beams.

Online

  • Nothing

L3

  • Still looking into the differences from old and new filter configuration and trace down where they loose events. The result will be presented next week and also in a BAD #1769.

Data Quality and Monitoring

  • Still looking into the pi0 yield change (up), pi0 mass change (unstable), pi0 width change (down), signal over background ratio change (up).

Trickle Injection Monitoring

  • Nothing

Trigger Configuration

AOB

  • Nothing

This page is created by Kim Hojeong (hojeong@slac); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.